OpenStudy (anonymous):

A box is has dimensions of 14 inches long, 1.5 feet wide, and 7 inches high. What is the volume of the box? The formula for the volume is V = l • w • h. (1 point) 1.02 cubic inches (in3) 12.25 cubic inches (in3) 147 cubic inches (in3) 1764 cubic inches (in3)

OpenStudy (anonymous):

\[v=14\times1.5\times7\] volume=147

OpenStudy (anonymous):

Basically, just use the Volume formula: Volume = Length*width*height Input the values and you get your volume. So: \[14 \times 1.5 \times 7 = ?\] Solve that to get your answer

OpenStudy (anonymous):

Not quite, you need to be carful with the units. Two of the lengths are in inches but one is in feet. You first need to convert 1.5 feet = 18 inches. THEN you can do: V=l x w x h V=14 x 18 x 7 V= 1764 cubic inches.

OpenStudy (anonymous):

If you want you can do it all in one step: \[V = l \times w \times h = 14" \times (1.5ft \times \frac{ 12" }{ 1ft }) \times 7" \]